in how many ways can a committee of three men and four women be formed from a group of 12 men and 12 women?

Answer by Edwin McCravy(20086)

in how many ways can a committee of three men and four women be formed from a group of 12 men and 12 women?

You can choose the three men in 12C3 ways, which is 220 ways.

For each of those 220 ways to chose three men, there are 12C4, which is

495 ways to choose the four women, so the answer is

12C3·12C4 = 330·495 = 108900 ways.

Edwin